OpenStudy (anonymous):

The vertex of this parabola is at (4,-3). When the x-value is 5, the y-value is at -6. What is the coefficient of the squared expression in the parabola's equation ? A. 3 B. -2 C. -3 D. 2

OpenStudy (anonymous):

Help me please :(

OpenStudy (sherribelle):

Use the formula: y = a(x - h)^2 + k vertex = (h,k)

OpenStudy (anonymous):

but how do i know which numbers to plug in to which variable ?

OpenStudy (sherribelle):

The vertex = (h,k) So y = -6 a = is the variable you're trying to find x = 5 h = 4 k = -3

OpenStudy (sherribelle):

So whatever the vertex is, is what you plug into h and k. It should look like: -6 = a(5-4)^2 - 3

OpenStudy (sherribelle):

The app Mathway is really helpful once you have everything set up.

OpenStudy (anonymous):

so i got 3 ?

OpenStudy (sherribelle):

Mathway got -3, let me work it out on paper.

OpenStudy (sherribelle):

I got 3 as well.